Suppose that you have a formula to measure the loudness of sounds on a logarithmic scale. You do not know what base to use for your logarithms, but you have determined by experiment that log(b)2=3.4 and log(b)3=5.4. Use this information and the properties of logarithms to calculate log(b)24.
thank in advance

Answer by stanbon(48535)

but you have determined by experiment that log(b)2=3.4 and log(b)3=5.4. Use this information and the properties of logarithms to calculate log(b)24.
--------------
----------
24=2^3*3
log(b)24 = log(b)[2^3*3]
=3log(b)2 +log(b)3
=3(3.4) + 5.4
=15.6
=============
Cheers,
Stan H.